OpenStudy (anonymous):

A mechanic charges $330 for 6 hours of labor. How much would she charge for 9 hours of labor? A. $495 B. $339 C. $336 D. $345

OpenStudy (anonymous):

B

OpenStudy (anonymous):

I think you are right do you know why?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

divide 330 into 6

OpenStudy (anonymous):

tell me what you get

OpenStudy (anonymous):

55

OpenStudy (anonymous):

so He charges 55 dollars an hour

OpenStudy (anonymous):

then 55 * 9 =339

OpenStudy (anonymous):

woops wrong srry

OpenStudy (jdoe0001):

\(\large \begin{array}{ccllll} \$&hours \\\hline\\ 330&6\\ x&9 \end{array}\implies \cfrac{330}{x}=\cfrac{6}{9}\)

OpenStudy (astrophysics):

You can make a ratio \[\frac{ 330 }{ x } = \frac{6 }{ 9 }\]

OpenStudy (anonymous):

you want to know how much he charges for 9 hours all you have to do is multiply the number of hours by how much is charged per hour and you will have your answer!

OpenStudy (anonymous):

495

OpenStudy (anonymous):

so its A

OpenStudy (astrophysics):

Yup

OpenStudy (anonymous):

Awesome!

OpenStudy (anonymous):

thankes